OpenStudy (anonymous):

mark is twice as old as lota. three years ago, the product of their ages is four more than twice the age of lota now. how old are they now?

OpenStudy (anonymous):

Let present age of mark be m and present age of lota be l. From first statement. m=2l. From second statement (m-3)(l-3) = 2l+4 Substitute for m in second equation from first equation. (2l-3)(l-3) = 2l+4 \[2l^{2}-9l+9 = 2l+4\] \[2l^{2}-11l+5 = 0\] Solving this equation will give you l = 1/2, 5. As age cannot be 1/2. Present age of lota is 5 and that of mark is 10 years